# Modeling Innovation Ecosystem Dynamics through Interacting Reinforced Bernoulli Processes

### v1

@@ Appendix A Definitions of novelty in the literature
    The most recent stream is the one of Natural Language Programming (NPL) and text analyses of patents (Abbas et al. 2014, Gerken and Moehrle 2012). In essence in Gerken and Moehrle (2012) the process of calculating patent novelty compares the text of a new patent with all previously filed patents. The similarity between the new patent and each prior one is measured, and the highest similarity score (the patent’s “oldness”) is subtracted from $1$ to determine the novelty score. Essentially, the more distinct a new patent is from earlier patents, the higher its novelty. This approach emphasizes the unique content of a patent by minimizing similarities with existing technologies. In Abbas et al. (2014), semantic analysis is mentioned as a developing method, with an emphasis on how linguistic analysis (e.g., analyzing keywords, abstracts) is becoming crucial in identifying novel patents. But, their definition of novelty primarily revolves around citation gaps [^14] If a patent doesn’t reference many previous patents (or none at all), it suggests that the new patent is introducing ideas not heavily reliant on earlier work. This is seen as a sign of novelty because it indicates that the patent isn’t just building on existing knowledge, but possibly presenting something new. and technological distance [^15] This refers to how different a patent is from those in similar fields. If a patent belongs to a new or less crowded technological class (based on how patents are categorized), it suggests that the patent is more novel because it explores a technological area that hasn’t been fully developed yet. (e.g., patents in new technological classes). More recently Kelly et al. (2021) defined, again, patent novelty by focusing on textual similarity. They measure novelty by comparing the text content of patents to both previous and future patents. A patent is considered novel if it is distinct from earlier patents but also relevant to future innovations. Their approach uses textual analysis to identify patents that contribute significantly to technological progress, and these are often predictive of future citations and market value.
@@
In the main text we adopt a category-version of the Squicciarini et al. (2013) index as our primary measure of patent novelty due to its widespread recognition and methodological robustness within the economics of innovation literature (Higham et al. 2021, Dehghani et al. 2023). This index, endorsed and implemented by the OECD in its patent quality indicators framework, captures the diversity of technological fields cited by a patent-interpreted as a proxy for knowledge recombination and inventive originality. It has also been applied across a range of recent empirical studies (Dehghani et al. 2023, Gao and Lazarova 2023, Angori et al. 2024) reflecting its growing standardization and comparability across datasets and time. Compared to newer text-based or machine-learning approaches, the Squicciarini et al. (2013)’s index offers a better replicability, and ease of modification in a cross-category setting as ours. Notice, however that our setting is general enough to allow for other measures of patent novelty as well.
@@ Appendix B Proofs of the main results and technical details for the statistical test
Proof.
Proof of Theorem 3.1 For each time step $t$, set ${{\mathbf{X}}}_{t}=(X_{t,1},\dots,X_{t,N})^{\top}$, ${{\mathbf{P}}}_{t}=(P_{t,1},\dots,P_{t,N})^{\top}$ and ${{\boldsymbol{\theta}}}=(\theta_{1},\dots,\theta_{N})^{\top}$. Hence, the vectorial dynamics for the random vector ${{\mathbf{P}}}_{t}$ is, for $t\geq 0$,
[…]
${{\mathbf{P}}}_{t+1}$ $=\Big{(}1-\frac{1}{t+1}\Big{)}{{\mathbf{P}}}_{t}+\frac{1}{t+1}\Gamma^{\top}{{\mathbf{P}}}_{t+1}+O(1/t^{2}){\boldsymbol{1}}$
$={{\mathbf{P}}}_{t}-\frac{1}{t+1}(I-\Gamma^{\top}){{\mathbf{P}}}_{t}+\frac{1}{t+1}\Gamma^{\top}\Delta{{\mathbf{M}}}^{\prime}_{t+1}+O(1/t^{2}){\boldsymbol{1}},$
@@
(B.2) $\zeta_{0}(x)=1,\qquad\zeta_{t}(x)=\frac{\Gamma(t+x)}{\Gamma(t)}\sim t^{x}\uparrow+\infty\,.$
@@
(B.3) $\frac{1}{\zeta_{t+1}(x)\zeta_{t+1}(1-x)}=$
@@
Hence, multiplying (B.1) by $\zeta_{t+1}(1-\gamma^{*})$ and using the relation
@@
${{\mathbf{A}}}_{t+1}$ $=\Big{(}1-\frac{1}{t+1}(I-\Gamma^{\top})\Big{)}\frac{\zeta_{t+1}(1-\gamma^{*})}{\zeta_{t}(1-\gamma^{*})}{{\mathbf{A}}}_{t}+\frac{1}{t+1}\Gamma^{\top}\Delta{{\mathbf{M}}}_{t+1}+O\Big{(}\frac{\zeta_{t+1}(1-\gamma^{*})}{t^{2}}\Big{)}{\boldsymbol{1}}$
$=\Big{(}1-\frac{1}{t+1}(I-\Gamma^{\top})\Big{)}\Big{(}1+\frac{1-\gamma^{*}}{t+1}+O(1/t^{2})\Big{)}{{\mathbf{A}}}_{t}+\frac{1}{t+1}\Gamma^{\top}\Delta{{\mathbf{M}}}_{t+1}+O\Big{(}\frac{\zeta_{t+1}(1-\gamma^{*})}{t^{2}}\Big{)}{\boldsymbol{1}}$
$={{\mathbf{A}}}_{t}-\frac{1}{t+1}(\gamma^{*}I-\Gamma^{\top}){{\mathbf{A}}}_{t}+\frac{1}{t+1}\Gamma^{\top}\Delta{{\mathbf{M}}}_{t+1}+O\Big{(}\frac{\zeta_{t+1}(1-\gamma^{*})}{t^{2}}\Big{)}{\boldsymbol{1}}\,,$
@@
Now, we are going to apply Theorem B.2 of Section B.1 with $\Phi=\Gamma$, $\phi^{*}=\gamma^{*}$ and ${\mathcal{F}}=({\mathcal{F}}_{t})_{t}$ the natural filtration associated to the model, i.e. ${\mathcal{F}}_{t}=\sigma(X_{n,h}:\,n\leq t,\,h=1,\dots,N)$. To this purpose, we choose ${\mathbf{u}}$ and ${\mathbf{v}}$ as the left and the right eigenvectors of $\Gamma$ associated to $\gamma^{*}$, with strictly positive components and such that ${\mathbf{v}}^{\top}{\boldsymbol{1}}=1$ and ${\mathbf{v}}^{\top}{\mathbf{u}}=1$. (Recall that $\gamma^{*}$ is simple and it is possible to choose the components of these vectors all strictly positive because of the Frobenius-Perron theory.) We set $\widetilde{A}_{t}={\mathbf{v}}^{\top}{\mathbf{A}}_{t}=\zeta_{t}(1-\gamma^{*}){\mathbf{v}}^{\top}{\mathbf{P}}_{t}$. First of all, we observe that assumptions (i) and (ii) of Theorem B.2 are satisfied because of ($\mathcal{A}1$) and ($\mathcal{A}2$), that imply $\gamma^{*}\in(0,1]$. Moreover, also assumption (iii) is verified: indeed, we have $\sup_{t}E[\widetilde{A}_{t}]<+\infty$ as $\widetilde{A}_{t}$ is non-negative (see Remark B.1). Finally, we have
@@
where $\widetilde{P}_{\infty}=\widetilde{A}_{\infty}$ is a square-integrable non-negative random variable, which is the almost sure limit of the above defined $\widetilde{A}_{t}$, or equivalently (by (B.2)), of $t^{1-\gamma^{*}}{\mathbf{v}}^{\top}{\mathbf{P}}_{t}$
It remains to prove that $P(\widetilde{P}_{\infty}>0)=1$. To this purpose, we use (Aletti et al. 2023a, Theorem S1.3). Indeed, setting $\widetilde{P}_{t}={\mathbf{v}}^{\top}{\mathbf{P}}_{t}$, we have
(B.4) $\widetilde{P}_{0}$ $={\mathbf{v}}^{\top}{\mathbf{P}}_{0}=\frac{1}{c}{\mathbf{v}}^{\top}\boldsymbol{\theta}>0$
@@
Hence, from (Aletti et al. 2023a, Theorem S1.3) applied to ${\mathcal{V}}=(\mathcal{V}_{t})$ with $\delta=\gamma^{*}$ (note that $E[Y_{t+1}|\mathcal{F}_{t}]=\gamma^{*}\widetilde{P}_{t}=\gamma^{*}\mathcal{V}_{t}+O(\ln(t)/t)$, we get that $t^{1-\gamma^{*}}\mathcal{V}_{t}$ converges almost surely to a strictly positive finite random variable. This random variable is obviously also the almost sure limit of $t^{1-\gamma^{*}}\widetilde{P}_{t}$ and so we can conclude that $P(\widetilde{P}_{\infty}>0)=1$. ∎∎
Proof.
Proof of Corollary 3.2 In the previous proof, we have proven that the limit random vector of $t^{1-\gamma^{*}}{\mathbf{P}}_{t}$ is ${\mathbf{P}}_{\infty}=\widetilde{P}_{\infty}{\mathbf{u}}$, where here $\widetilde{P}_{\infty}$ refers to a precise choice for the vector ${\mathbf{u}}$, that is $\widetilde{P}_{\infty}=\widetilde{P}_{\infty}({\mathbf{u}})$. If we choose a different (left) eigenvector ${\mathbf{u}}^{\prime}$ of $\Gamma$ associated to $\gamma^{*}$, then we necessarily have have ${\mathbf{u}}^{\prime}=C{\mathbf{u}}$ with $C\neq 0$ and so we can write the limit random vector as ${\mathbf{P}}_{\infty}=\widetilde{P}_{\infty}({\mathbf{u}}^{\prime}){\mathbf{u}}^{\prime}$ with $\widetilde{P}_{\infty}({\mathbf{u}}^{\prime})=\widetilde{P}_{\infty}({\mathbf{u}})/C$. Summing up, for any choice of the (left) eigenvector ${\mathbf{u}}$, the limit random vector can be written as ${\mathbf{P}}_{\infty}=\widetilde{P}_{\infty}({\mathbf{u}}){\mathbf{u}}$ with a suitable square-integrable random variable $\widetilde{P}_{\infty}({\mathbf{u}})$ such that $P(\widetilde{P}_{\infty}({\mathbf{u}})\neq 0)=1$ and $\widetilde{P}_{\infty}(C{\mathbf{u}})=\widetilde{P}_{\infty}({\mathbf{u}})/C$. In other words, each random variable $P_{\infty,h}$ can be factorized in the product of a deterministic term specific for each $h$, i.e. $u_{h}$, and a common random term, i.e. $\widetilde{P}_{\infty}({\mathbf{u}})$.
@@
$t^{1-\gamma^{*}}Var[X_{t+1,h}]=t^{1-\gamma^{*}}E[P_{t,h}(1-P_{t,h})]\to\begin{cases}E[P_{\infty,h}]=u_{h}E[\widetilde{P}_{\infty}({\mathbf{u}})]\;&\text{for}\gamma^{*}<1\\ E[P_{\infty,h}(1-P_{\infty,h})]=u_{h}E\left[\widetilde{P}_{\infty}({\mathbf{u}})(1-u_{h}\widetilde{P}_{\infty}({\mathbf{u}}))\right]\;&\text{for }\gamma^{*}=1\,.\end{cases}$
Therefore, since the sign of the components of ${\mathbf{u}}$ (which is the same for all of them) necessarily coincides with the one of $\widetilde{P}_{\infty}({\mathbf{u}})$, in order to obtain the first limit relation in Corollary 3.2, we can set $\alpha({\mathbf{u}})$ equal to $|E[\widetilde{P}_{\infty}(u)]|$ when $\gamma^{*}<1$ and equal to $|E\left[\widetilde{P}_{\infty}({\mathbf{u}})(1-u_{h}\widetilde{P}_{\infty}({\mathbf{u}}))\right]|$ when $\gamma^{*}=1$, so that the two possible cases in the above formula can be summarized as $|u_{h}|\alpha({\mathbf{u}})$. Moreover, for each pair $h\neq j$, since $X_{t+1,h}$ and $X_{t+1,j}$ are conditionally independent given ${\mathcal{F}}_{t}$, we get
@@
Proof.
Proof of Theorem 3.3 From Theorem 3.1, using the same notation adopted in its proof, we get
@@
(B.5) $S_{t,h}\stackrel{{\scriptstyle a.s.}}{{\sim}}S_{\infty,h}\,t^{\gamma^{*}}~{}\mbox{ with }S_{\infty,h}=\frac{P_{\infty,h}}{\gamma^{*}}\,.$
(We can also note that, similarly, since the convergence in Theorem 3.1 is also in mean, we also have $E[S_{t,h}]=\sum_{n=0}^{t}E[X_{n,h}]=\sum_{n=0}^{t}E[P_{n,h}]\sim t^{\gamma^{*}}E[S_{\infty,h}]$.) As a consequence of (B.5) and the fact that $P_{\infty,h}/P_{\infty,j}=u_{h}/u_{j}$, we obtain
@@
Proof.
Proof of Theorem 3.4 From Theorem 3.1 and Theorem 3.3, using the same notation adopted in their proofs, we have by ($\mathcal{A}3$)
@@
and so it is enough to apply (Williams 1991, sec. 12.15) in order to obtain $S_{t,k,h}=\sum_{n=1}^{t}X_{n,h}Y_{n,k}\stackrel{{\scriptstyle a.s.}}{{\sim}}\pi_{k}S_{\infty,h}\,t^{\gamma^{*}}$. ∎∎
@@
We here use the same choice of the eigenvectors ${\mathbf{v}}$ and ${\mathbf{u}}$ and the same notation adopted in the proof of Theorem 3.1. We recall that the dynamics of the above defined process ${\mathbf{A}}_{t}=\zeta_{t}(1-\gamma^{*}){\mathbf{P}}_{t}$ with $\zeta_{t}(\cdot)$ defined in (B.2), is
@@
${{\mathbf{A}}}_{t+1}$ $={{\mathbf{A}}}_{t}-\frac{1}{t+1}(\gamma^{*}I-\Gamma^{\top}){{\mathbf{A}}}_{t}+\frac{1}{t+1}\Gamma^{\top}\Delta{{\mathbf{M}}}_{t+1}+O\left(\frac{\zeta_{t+1}(1-\gamma^{*})}{t^{2}}\right){\boldsymbol{1}}\,,$

### v2

@@ Appendix A Definitions of novelty in the literature
    The most recent stream is the one of Natural Language Programming (NPL) and text analyses of patents (Abbas, Zhang and Khan 2014, Gerken and Moehrle 2012). In essence in Gerken and Moehrle (2012) the process of calculating patent novelty compares the text of a new patent with all previously filed patents. The similarity between the new patent and each prior one is measured, and the highest similarity score (the patent’s “oldness”) is subtracted from $1$ to determine the novelty score. Essentially, the more distinct a new patent is from earlier patents, the higher its novelty. This approach emphasizes the unique content of a patent by minimizing similarities with existing technologies. In Abbas, Zhang and Khan (2014), semantic analysis is mentioned as a developing method, with an emphasis on how linguistic analysis (e.g., analyzing keywords, abstracts) is becoming crucial in identifying novel patents. But, their definition of novelty primarily revolves around citation gaps [^12] If a patent doesn’t reference many previous patents (or none at all), it suggests that the new patent is introducing ideas not heavily reliant on earlier work. This is seen as a sign of novelty because it indicates that the patent isn’t just building on existing knowledge, but possibly presenting something new. and technological distance [^13] This refers to how different a patent is from those in similar fields. If a patent belongs to a new or less crowded technological class (based on how patents are categorized), it suggests that the patent is more novel because it explores a technological area that hasn’t been fully developed yet. (e.g., patents in new technological classes). More recently Kelly et al. (2021) defined, again, patent novelty by focusing on textual similarity. They measure novelty by comparing the text content of patents to both previous and future patents. A patent is considered novel if it is distinct from earlier patents but also relevant to future innovations. Their approach uses textual analysis to identify patents that contribute significantly to technological progress, and these are often predictive of future citations and market value.
@@
In the main text we adopt a category-version of the Squicciarini, Dernis and Criscuolo (2013) index as our primary measure of patent novelty due to its widespread recognition and methodological robustness within the economics of innovation literature (Higham, De Rassenfosse and Jaffe 2021, Dehghani et al. 2023). This index, endorsed and implemented by the OECD in its patent quality indicators framework, captures the diversity of technological fields cited by a patent-interpreted as a proxy for knowledge recombination and inventive originality. It has also been applied across a range of recent empirical studies (Dehghani et al. 2023, Gao and Lazarova 2023, Angori et al. 2024) reflecting its growing standardization and comparability across datasets and time. Compared to newer text-based or machine-learning approaches, the Squicciarini, Dernis and Criscuolo (2013)’s index offers a better replicability, and ease of modification in a cross-category setting as ours. Notice, however that our setting is general enough to allow for other measures of patent novelty as well.
@@ Appendix B Proofs of the main results and technical details for the statistical test
Proof of Theorem 3.1.
For each time step $t$, set ${{\mathbf{X}}}_{t}=(X_{t,1},\dots,X_{t,N})^{\top}$, ${{\mathbf{P}}}_{t}=(P_{t,1},\dots,P_{t,N})^{\top}$ and ${{\boldsymbol{\theta}}}=(\theta_{1},\dots,\theta_{N})^{\top}$. Hence, the vectorial dynamics for the random vector ${{\mathbf{P}}}_{t}$ is, for $t\geq 0$,
${{\mathbf{P}}}_{0}$ $={\boldsymbol{\theta}}/c\neq{\mathbf{0}}\quad\text{a.s.}$ (4)
${{\mathbf{P}}}_{t+1}$ $=\Big(1-\frac{1}{t+1}\Big){{\mathbf{P}}}_{t}+\frac{1}{t+1}\Gamma^{\top}{{\mathbf{P}}}_{t+1}+O(1/t^{2}){\mathbf{1}}$
$={{\mathbf{P}}}_{t}-\frac{1}{t+1}(I-\Gamma^{\top}){{\mathbf{P}}}_{t}+\frac{1}{t+1}\Gamma^{\top}\Delta{{\mathbf{M}}}^{\prime}_{t+1}+O(1/t^{2}){\mathbf{1}},$
@@
$\zeta_{0}(x)=1,\qquad\zeta_{t}(x)=\frac{\Gamma(t+x)}{\Gamma(t)}\sim t^{x}\uparrow+\infty\,.$ (5)
@@
$\frac{1}{\zeta_{t+1}(x)\zeta_{t+1}(1-x)}=$ (6)
@@
Hence, multiplying (4) by $\zeta_{t+1}(1-\gamma^{*})$ and using the relation
@@
${{\mathbf{A}}}_{t+1}$ $=\Big(1-\frac{1}{t+1}(I-\Gamma^{\top})\Big)\frac{\zeta_{t+1}(1-\gamma^{*})}{\zeta_{t}(1-\gamma^{*})}{{\mathbf{A}}}_{t}+\frac{1}{t+1}\Gamma^{\top}\Delta{{\mathbf{M}}}_{t+1}+O\Big(\frac{\zeta_{t+1}(1-\gamma^{*})}{t^{2}}\Big){\mathbf{1}}$
$=\Big(1-\frac{1}{t+1}(I-\Gamma^{\top})\Big)\Big(1+\frac{1-\gamma^{*}}{t+1}+O(1/t^{2})\Big){{\mathbf{A}}}_{t}+\frac{1}{t+1}\Gamma^{\top}\Delta{{\mathbf{M}}}_{t+1}+O\Big(\frac{\zeta_{t+1}(1-\gamma^{*})}{t^{2}}\Big){\mathbf{1}}$
$={{\mathbf{A}}}_{t}-\frac{1}{t+1}(\gamma^{*}I-\Gamma^{\top}){{\mathbf{A}}}_{t}+\frac{1}{t+1}\Gamma^{\top}\Delta{{\mathbf{M}}}_{t+1}+O\Big(\frac{\zeta_{t+1}(1-\gamma^{*})}{t^{2}}\Big){\mathbf{1}}\,$,
@@
Now, we are going to apply Theorem B.2 of Section B.1 with $\Phi=\Gamma$, $\phi^{*}=\gamma^{*}$ and ${\mathcal{F}}=({\mathcal{F}}_{t})_{t}$ the natural filtration associated to the model, i.e. ${\mathcal{F}}_{t}=\sigma(X_{n,h}:\,n\leq t,\,h=1,\dots,N)$. To this purpose, we choose $\mathbf{u}$} and $\mathbf{v}$} as the left and the right eigenvectors of $\Gamma$ associated to $\gamma^{*}$, with strictly positive components and such that ${\mathbf{v}}^{\top}{\mathbf{1}}=1$ and ${\mathbf{v}}^{\top}{\mathbf{u}}=1$. (Recall that $\gamma^{*}$ is simple and it is possible to choose the components of these vectors all strictly positive because of the Frobenius-Perron theory.) We set $\widetilde{A}_{t}={\mathbf{v}}^{\top}{\mathbf{A}}_{t}=\zeta_{t}(1-\gamma^{*}){\mathbf{v}}^{\top}{\mathbf{P}}_{t}$. First of all, we observe that assumptions (i) and (ii) of Theorem B.2 are satisfied because of ($\mathcal{A}1$) and ($\mathcal{A}2$), that imply $\gamma^{*}\in(0,1]$. Moreover, also assumption (iii) is verified: indeed, we have $\sup_{t}E[\widetilde{A}_{t}]<+\infty$ as $\widetilde{A}_{t}$ is non-negative (see Remark 2). Finally, we have
@@
where $\widetilde{P}_{\infty}=\widetilde{A}_{\infty}$ is a square-integrable non-negative random variable, which is the almost sure limit of the above defined $\widetilde{A}_{t}$, or equivalently (by (5)), of $t^{1-\gamma^{*}}{\mathbf{v}}^{\top}{\mathbf{P}}_{t}$ .
It remains to prove that $P(\widetilde{P}_{\infty}>0)=1$. To this purpose, we use (Aletti, Crimaldi and Ghiglietti 2023, Theorem S1.3). Indeed, setting $\widetilde{P}_{t}={\mathbf{v}}^{\top}{\mathbf{P}}_{t}$, we have
$\widetilde{P}_{0}$ $={\mathbf{v}}^{\top}{\mathbf{P}}_{0}=\frac{1}{c}{\mathbf{v}}^{\top}\boldsymbol{\theta}>0$ (7)
@@
Hence, from (Aletti, Crimaldi and Ghiglietti 2023, Theorem S1.3) applied to ${\mathcal{V}}=(\mathcal{V}_{t})$ with $\delta=\gamma^{*}$ (note that $E[Y_{t+1}|\mathcal{F}_{t}]=\gamma^{*}\widetilde{P}_{t}=\gamma^{*}\mathcal{V}_{t}+O(\ln(t)/t)$, we get that $t^{1-\gamma^{*}}\mathcal{V}_{t}$ converges almost surely to a strictly positive finite random variable. This random variable is obviously also the almost sure limit of $t^{1-\gamma^{*}}\widetilde{P}_{t}$ and so we can conclude that $P(\widetilde{P}_{\infty}>0)=1$. ∎
Proof of Corollary 3.1.1.
In the previous proof, we have proven that the limit random vector of $t^{1-\gamma^{*}}{\mathbf{P}}_{t}$ is ${\mathbf{P}}_{\infty}=\widetilde{P}_{\infty}{\mathbf{u}}$, where here $\widetilde{P}_{\infty}$ refers to a precise choice for the vector $\mathbf{u}$}, that is $\widetilde{P}_{\infty}=\widetilde{P}_{\infty}({\mathbf{u}})$. If we choose a different (left) eigenvector $\mathbf{u}$$^{\prime}$ of $\Gamma$ associated to $\gamma^{*}$, then we necessarily have have ${\mathbf{u}}^{\prime}=C{\mathbf{u}}$ with $C\neq 0$ and so we can write the limit random vector as ${\mathbf{P}}_{\infty}=\widetilde{P}_{\infty}({\mathbf{u}}^{\prime}){\mathbf{u}}^{\prime}$ with $\widetilde{P}_{\infty}({\mathbf{u}}^{\prime})=\widetilde{P}_{\infty}({\mathbf{u}})/C$. Summing up, for any choice of the (left) eigenvector $\mathbf{u}$}, the limit random vector can be written as ${\mathbf{P}}_{\infty}=\widetilde{P}_{\infty}({\mathbf{u}}){\mathbf{u}}$ with a suitable square-integrable random variable $\widetilde{P}_{\infty}({\mathbf{u}})$ such that $P(\widetilde{P}_{\infty}({\mathbf{u}})\neq 0)=1$ and $\widetilde{P}_{\infty}(C{\mathbf{u}})=\widetilde{P}_{\infty}({\mathbf{u}})/C$. In other words, each random variable $P_{\infty,h}$ can be factorized in the product of a deterministic term specific for each $h$, i.e. $u_{h}$, and a common random term, i.e. $\widetilde{P}_{\infty}({\mathbf{u}})$.
@@
$t^{1-\gamma^{*}}Var[X_{t+1,h}]=t^{1-\gamma^{*}}E[P_{t,h}(1-P_{t,h})]\\ \longrightarrow\begin{cases}E[P_{\infty,h}]=u_{h}E[\widetilde{P}_{\infty}({\mathbf{u}})]\;&\text{for }\gamma^{*}<1\\ E[P_{\infty,h}(1-P_{\infty,h})]=u_{h}E\left[\widetilde{P}_{\infty}({\mathbf{u}})(1-u_{h}\widetilde{P}_{\infty}({\mathbf{u}}))\right]\;&\text{for }\gamma^{*}=1\,.\end{cases}$
Therefore, since the sign of the components of $\mathbf{u}$} (which is the same for all of them) necessarily coincides with the one of $\widetilde{P}_{\infty}({\mathbf{u}})$, in order to obtain the first limit relation in Corollary 3.1.1, we can set $\alpha({\mathbf{u}})$ equal to $|E[\widetilde{P}_{\infty}(u)]|$ when $\gamma^{*}<1$ and equal to $|E\left[\widetilde{P}_{\infty}({\mathbf{u}})(1-u_{h}\widetilde{P}_{\infty}({\mathbf{u}}))\right]|$ when $\gamma^{*}=1$, so that the two possible cases in the above formula can be summarized as $|u_{h}|\alpha({\mathbf{u}})$. Moreover, for each pair $h\neq j$, since $X_{t+1,h}$ and $X_{t+1,j}$ are conditionally independent given ${\mathcal{F}}_{t}$, we get
@@
Proof of Theorem 3.2.
From Theorem 3.1, using the same notation adopted in its proof, we get
@@
$S_{t,h}\stackrel{{\scriptstyle a.s.}}{{\sim}}S_{\infty,h}\,t^{\gamma^{*}}~\mbox{ with }S_{\infty,h}=\frac{P_{\infty,h}}{\gamma^{*}}\,.$ (8)
(We can also note that, similarly, since the convergence in Theorem 3.1 is also in mean, we also have $E[S_{t,h}]=\sum_{n=0}^{t}E[X_{n,h}]=\sum_{n=0}^{t}E[P_{n,h}]\sim t^{\gamma^{*}}E[S_{\infty,h}]$.) As a consequence of (8) and the fact that $P_{\infty,h}/P_{\infty,j}=u_{h}/u_{j}$, we obtain
@@
Proof of Theorem 3.3.
From Theorem 3.1 and Theorem 3.2, using the same notation adopted in their proofs, we have by ($\mathcal{A}3$)
@@
and so it is enough to apply (Williams 1991, sec. 12.15) in order to obtain $S_{t,k,h}=\sum_{n=1}^{t}X_{n,h}Y_{n,k}\stackrel{{\scriptstyle a.s.}}{{\sim}}\pi_{k}S_{\infty,h}\,t^{\gamma^{*}}$. ∎
@@
We here use the same choice of the eigenvectors $\mathbf{v}$} and $\mathbf{u}$} and the same notation adopted in the proof of Theorem 3.1. We recall that the dynamics of the above defined process ${\mathbf{A}}_{t}=\zeta_{t}(1-\gamma^{*}){\mathbf{P}}_{t}$ with $\zeta_{t}(\cdot)$ defined in (5), is
@@
${{\mathbf{A}}}_{t+1}$ $={{\mathbf{A}}}_{t}-\frac{1}{t+1}(\gamma^{*}I-\Gamma^{\top}){{\mathbf{A}}}_{t}+\frac{1}{t+1}\Gamma^{\top}\Delta{{\mathbf{M}}}_{t+1}+O\left(\frac{\zeta_{t+1}(1-\gamma^{*})}{t^{2}}\right){\mathbf{1}}\,$,